# Decision by Supervised Learning with Deep Ensembles: A Practical Framework for Robust Portfolio Optimization

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2025submissionid: 106ccs: Applied computing Computers in other domainsccs: Mathematics of computing Bayesian computationccs: Computing methodologies Ensemble methods
Figure 1. Graphical illustrations of Machine Learning-based Portfolio Optimization methods
[…]
This is a graphical illustrations of Prediction-Focused Learning, Decision-Focused Learning, End-to-End Learning and Decision by Supervised Learning
@@ 3.1. Preliminaries: Prediction-Focused Learning, Decision-Focused Learning, and End-to-End Learning
[…]
@@
[…]
@@
[…]
@@ 4.3. Backtesting Experiment (RQ1, RQ2)
Figure 2. Comparing cumulative return trajectory with classical portfolio strategies.
[…]
We compare the backtesting results with classical portfolio strategies, such as equal-weighted, value-weighted portfolios and mSSRM. We visualize cumulative return of different strategies for each universe.
@@ 4.4. Ensemble Effect Analysis (RQ3)
Figure 3. Impact of Ensemble Size on Performance
[…]
Ensemble effect experiment on NASDAQ 100 Rolling universe and S&P500 Rolling universe
@@ Acknowledgements.
This document was prepared by the AI Solution Division of Mirae Asset Global Investments for informational purposes only. Mirae Asset Global Investments makes no warranty as to the accuracy, completeness, or reliability of the information contained herein and accepts no legal responsibility for any consequences arising from its use. This document does not constitute investment advice, a recommendation, or an offer to buy or sell any financial products, and should not be used as a basis for investment decisions.

### v6

@@ Abstract.
copyright: acmlicensedjournalyear: 2025conference: The 34th ACM International Conference on Information and Knowledge Management; November 10–-14, 2025; Seoulsubmissionid: 1ccs: Applied computing Computers in other domainsccs: Mathematics of computing Bayesian computationccs: Computing methodologies Ensemble methods
Figure 1. Graphical illustrations of Machine Learning-based Portfolio Optimization methods:
@@ 3.1. Preliminaries: Prediction-Focused Learning, Decision-Focused Learning, and End-to-End Learning
(1) $\text{Stage 1 objective: }\hat{\theta}=\operatorname*{arg\,min}_{\theta}\mathcal{L}(\hat{y},y),\quad\hat{y}=f_{\theta}(x)$
@@
(3) $\text{Objective: }\hat{\theta}=\operatorname*{arg\,max}_{\theta}\mathcal{P}(w^{*}(\hat{y}),y),\quad\hat{y}=f_{\theta}(x)$
@@
(4) $\text{Objective: }\hat{\theta}=\operatorname*{arg\,max}_{\theta}\mathcal{P}(\hat{w},y),\quad\hat{w}=f_{\theta}(x)$
@@ 4.3. Backtesting Experiment (RQ1, RQ2)
Figure 2. Comparing cumulative return trajectory with classical portfolio strategies.:
@@ 4.4. Ensemble Effect Analysis (RQ3)
Figure 3. Impact of Ensemble Size on Performance:
@@ Acknowledgements.
This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all. This document was prepared by the AI Solution Team at Mirae Asset Global Investments for informational purposes only. Mirae Asset Global Investments makes no representation or warranty as to the accuracy, completeness, or reliability of the information contained herein and accepts no liability for any consequences arising from its use. This document does not constitute investment advice, a recommendation, or an offer to buy or sell any financial products, and should not be relied upon as a basis for investment decisions.